# Quantum Optical Electron Pulse Shaper

### v1

@@ IV.3 Short Pulse Gating
Until now, we have focused on modulation using long optical pulses that fully encompass the electron pulse. However, if the electron current is sufficiently high, it becomes possible to use optical pulses that are several times shorter than the electron pulse. This method requires shorter propagation distances for compression and is compatible with electron pulses of higher central energies. Additionally, it allows the final compressed sideband duration to be reduced below the photoemission duration, making the approach suitable even for pulses with long photoemission times.
We model the short pulse gating on an electron wavepacket with increased central kinetic energy of $30\,\mathrm{keV}$. The energy bandwidth $\delta E$ and the coherence time $\tau_{\text{coh}}$ remain unchanged. The initial duration of the electron pulse is increased to $\tau_{0}=50\,\mathrm{fs}$ [Fig 4(a)] and the duration after dispersive propagation remains $\tau_{\text{el}}=250\,\mathrm{fs}$ FWHM. The GDD coefficient is $\alpha=321\,\mathrm{fs^{2}}$, and the TDG coefficient is $\beta=0$ for simplicity.
While maintaining a constant spectral width of the optical pulse of $0.5\,\mathrm{eV}$, we set the GDD coefficient to $\varphi^{\prime\prime}=-\alpha/n$, where $n=11$. This is consistent with increasing the chirp $a$ while decreasing the pulse duration $\tau_{\text{opt}}$ approximately $n$-times. After the interaction a short, steeply chirped sideband centered around $1.2\,\mathrm{eV}$ is produced [Fig 4(b)]. To optimize the population of the first side-band we set $g=1$
The propagation distance required for the compression of the first sideband is then reduced to $z_{\text{com}}\approx\frac{m_{e}v^{2}\alpha}{\hbar(n-1)}$ while the time duration of the pulse after compression is reduced down to $\tau_{\text{com}}\approx\tau_{\text{0}}/(n-1)$. Specifically for $n=11$, where $4\%$ of the electrons from the initial pulse are transferred to the first sideband, the compression distance is $z_{\text{com}}\approx 32\,\mathrm{cm}$ and the duration of the compressed electron pulse is $\tau_{\text{com}}\approx 6\,\mathrm{fs}$ [Fig 4(c,d)].
@@
Figure 4: (a) Spectrogram of the long initial pulse with FWHM duration 50 fs. Spectrograms (b) of the chirped electron pulse after interaction with a short optical pulse, (c) of the electron pulse state after propagation in free space for $\approx$ 30 cm, (d) close up of the compressed sideband with FWHM temporal duration of $\approx$ 10 fs. Spectrograms (e) of the electron pulse state after interaction with phase- and amplitude-modulated optical field, (f) after propagation for 21 cm, (g) close up of the compressed pulse train.
IV.4 Periodic Gating
@@
The proposed scheme can however be used to produce tailored shape in the sideband, that will compress into a train of few-fs pulses in the compression point. While a single few-femtosecond laser pulse generates a single ultrashort electron sub-pulse, a longer optical pulse with periodic modulation in phase and amplitude can produce a train of ultrashort electron sub-pulses.
We start with 30 keV, 50 fs long electron pulse, stretched to 250 fs FWHM, as in the previous subsection. The optical pulse properties for this case can be introduced clearly in the time domain via the time-dependent frequency $\omega(t)$. As we state above, periodic frequency modulation is required
$\omega(t)=\omega_{0}\left[1+m\cos{\left(\Omega t\right)}\right]+2at,$ (18)
@@
where $0<m\leq 1$ is the modulation depth and $\Omega$ is the modulation frequency, which should be a few times smaller than $\omega_{0}$ for this application. The chirp term is present to compensate for the zero-loss band chirp. If an optical field with cosine frequency modulation is used, the spacing between the sidebands is also periodically modulated. The optical field in the frequency domain can be written as
~$\begin{split}\tilde{\mathcal{E}}(\omega)\propto\sum_{k=-1}^{k=1}\sum_{n=-\infty}^{\infty}(1/2)^{|k|}i^{n}J_{n}\left(\frac{m\omega_{0}}{\Omega}\right)\times\\ \exp\left[-\frac{(\omega-\omega_{0,kn})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0,kn})^{2}}{2}\right],\end{split}$ (19)
@@
Observing the first side-band, we note that every half period of the frequency oscillation the electrons have negative chirp and then in the next half period the electrons have positive chirp. The chirp can be precisely controlled by the modulation depth $m$ and frequency $\Omega$. In our case we choose $\Omega=\omega/20$ and $m=0.25$.
@@
Ideally, only the parts of the sideband which have negative chirp must be populated. This is achieved by amplitude modulation of the optical field $\sqrt{I(t)}\propto\left[1+\cos{\left(\Omega t\right)}\right]/2$. Now only within the selected half-periods is the light field strong enough for a side-band to be generated [Fig. 4(e)]. This leads to the generation of negatively chirped sub-pulses, separated in time by $1/2\pi\Omega$, and with chirp coefficients $\mathrm{d}E/\mathrm{d}t=-\hbar m\omega_{0}\Omega$. The compression occurs at $z_{\text{com}}\approx\frac{m_{e}v^{2}\alpha}{\hbar(m\omega_{0}\Omega/2a-1)}\approx 21\,\mathrm{cm}$ with the duration of the compressed pulse of $\tau_{\text{com}}\approx 6\,\mathrm{fs}$ [Fig. 4(f,g)]. The structured side-band contains around $10\%$ of the initial electron distribution.
@@
In the ideal case, a train of femtosecond pulses compressed several times below the initial duration of the pulse is generated. Considering realistic limitations, such as the finite bandwidth and the finite resolution of a standard light modulator, this application is still feasible, although minor distortions are to be expected. The temporal envelope of the side-band is expected to be shortened and the sideband modulation depth might be reduced (see Supplemental Material for details [51]).
V Discussion
The presented scheme is practically limited by three main factors. Firstly, the spectral and temporal properties of the optical pulses that can be realistically achieved with current shaping technologies. Secondly, we need to consider the phase-matching conditions of the chosen light-electron interaction over the broad spectrum of the optical pulse. Lastly, the available trajectory length within the electron microscope column determines the central energy of the electrons that will be used and the specific shape of the optical pulse.
@@ V Discussion
State-of-the-art light shapers enable precise control over the spectral phase of laser pulses. Our method requires broadband optical pulses with an energy bandwidth of about 0.5 eV to allow straightforward shaping through the introduction of GDD and third-order dispersion TOD. More complex pulse shaping—such as that needed for generating periodic pulse trains—demands a light shaper with both high spectral resolution and sufficient bandwidth. For instance, a shaper with 300 pixels and a bandwidth of 0.7 eV centered around 1.2 eV is expected to produce a pulse closely matching the ideal case. For further details, see Supplementary Material.
@@
In the main text, we approximate the coupling parameter $g(\omega)$ as constant over the bandwidth of the light pulse. In reality, the spectral dependence of $g(\omega)$ must be taken into account, and it may be necessary to adjust $I(\omega)$ to achieve the desired temporal profile of the generated sidebands. Furthermore, we do not address the effects of transverse momentum transfer during the interaction, which are expected to introduce additional nonlinear chirp due to the unequal trajectory lengths associated with different energies within the sidebands. However, we note that this effect can be mitigated by iteratively shaping the optical field to also compensate for electron delays arising from transverse effects.
Additional temporal broadening can arise from the fact that, in a focused beam, some electrons must traverse longer paths. This effect can be compensated by slightly defocusing the optical field at the membrane in the interaction plane. The phase curvature of the focused beam can then be used to counteract the trajectory length differences of the electrons.
VI Conclusion
@@
We theoretically demonstrate a technique for versatile free-electron manipulation through quantum-coherent interaction with intensity- and phase-shaped optical fields. This approach effectively transfers the state of the art of light pulse shaping techniques to the pulsed electron case, enabling precise control over the electron time-energy characteristics—including spectral width, temporal duration, and both second and higher-order dispersion—in a broadly applicable manner. We demonstrate that the quantum optical electron pulse shaper may achieve electron pulse compression of electrons to few-femtosecond pulse durations without additional spectral broadening of electrons, making this approach ideally suited for ultrafast experiments requiring high time, spatial and spectral resolutions.
Acknowledgements.
@@
Data Availability
The data supporting the findings of this study are openly available at [52].
@@
July 21, 2025
I Feasibility considerations
@@
~$\tilde{\mathcal{E}}(\omega)\propto\exp\left[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0})^{2}}{2}-i\frac{\varphi^{\prime\prime\prime}(\omega-\omega_{0})^{3}}{6}\right],$ (S1)
~$\begin{split}\tilde{\mathcal{E}}(\omega)\propto\sum_{k=-1}^{k=1}\sum_{n=-\infty}^{\infty}(1/2)^{|k|}i^{n}J_{n}\left(\frac{m\omega_{0}}{\Omega}\right)\times\exp\left[-\frac{(\omega-\omega_{0,kn})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0,kn})^{2}}{2}\right]\end{split}$ (S2)
@@
To model the ”realistic” light pulse, we first limit the spectral width by a rectangular window, which is non-zero for the interval $(\omega_{0}-\Delta\omega/2;\omega_{0}+\Delta\omega/2)$. Then we divide this interval by the number of pixels $N$ and assign a discrete amplitude and phase value to each of the frequencies, effectively truncating and down-sampling the light pulse frequency spectrum and spectral phase. Formally, the integration is replaced by the discrete sum
@@
On the other hand, in the case of the periodic frequency and intensity modulation of the optical pulse, the spectrum is more complicated, consisting of many peaks (see Fig. S3, blue line). While the discretization to 100 pixels is not detrimental to the peak resolution, important features of the spectrum are not transmitted through the spectral window, resulting in changes in the spectrum of the optical pulse Fig. S3, orange line. After interaction with the realistic optical pulse intensity spectrum, the sub-pulses have narrower spectrum [see Fig. (S4)]. However, our data demonstrate that downsampling of the spectral intensity and phase of the optical pulse shaper are not detrimental for the operation of the proposed electron pulse shaper.

### v2

@@ Abstract
Short Pulse Gating—Until now, we have focused on modulation using long optical pulses that fully encompass the electron pulse. However, if the electron current is sufficiently high, it becomes possible to use optical pulses that are several times shorter than the electron pulse. This method requires shorter propagation distances for compression and is compatible with electron pulses of higher central energies, that would require distances of several meters to compress with the previous sideband generation method. Additionally, it allows the final compressed sideband duration to be reduced below the photoemission duration, making the approach suitable even for pulses with longer photoemission times.
@@
Figure 4: (a) Spectrogram of the long initial pulse with FWHM duration 50 fs, (b) of the chirped electron pulse after interaction with a short optical pulse, (c) of the electron pulse state after propagation in free space for $\approx$ 32 cm, (d) close up of the compressed sideband with FWHM temporal duration of $\approx$ 10 fs. Spectrograms (e) of the electron pulse state after interaction with phase- and amplitude-modulated optical field, (f) after propagation for 21 cm, (g) close up of the compressed pulse train.
We model the short pulse gating on an electron wavepacket accelerated to central kinetic energy of $30\,\mathrm{keV}$ ($v_{0}=0.3c$). The energy bandwidth $\delta E$ remains unchanged. The initial duration of the electron pulse is increased to $\tau_{0}=50\,\mathrm{fs}$ [Fig. 4(a)] and the duration after acceleration and dispersive propagation remains $\tau_{\text{el}}=250\,\mathrm{fs}$ FWHM. The GDD coefficient is $\alpha=321\,\mathrm{fs^{2}}$, and the TDG coefficient is $\beta=0$ for simplicity.
While maintaining a constant spectral width of the optical pulse of $0.5\,\mathrm{eV}$, we set the GDD coefficient to $\varphi^{\prime\prime}=-\alpha/n$, where $n=11$. After the interaction a short, steeply chirped sideband centered around $1.2\,\mathrm{eV}$ is produced [Fig. 4(b)]. To optimize the population of the first side-band we set $g=1$.
The propagation distance required for the compression of the first sideband is then reduced to $z_{\text{com}}\approx\frac{m_{e}v_{\hbar\omega_{0}}^{2}\alpha}{\hbar(n-1)}$, with $v_{\hbar\omega_{0}}=v(E_{0}+\hbar\omega_{0})$, while the time duration of the pulse after compression is reduced down to $\tau_{\text{com}}\approx\tau_{\text{0}}/(n-1)$. Specifically for $n=11$, where $4\%$ of the electrons from the initial pulse are transferred to the first sideband, the compression distance is $z_{\text{com}}\approx 32\,\mathrm{cm}$ and the duration of the compressed electron pulse is $\tau_{\text{com}}\approx 6\,\mathrm{fs}$ [Fig. 4(c,d)].
@@
Periodic Gating—The proposed scheme can be used to produce tailored shape in the sideband, that will compress into a train of few-fs pulses in the compression point. While a single few-femtosecond laser pulse generates a single ultrashort electron sub-pulse, a longer optical pulse with periodic modulation in phase and amplitude can produce a train of ultrashort electron sub-pulses.
We start with the same electron pulse parameters as in the previous subsection. Regarding the laser field, periodic frequency modulation is applied
$\omega(\tau)=\omega_{0}\left[1+m\cos{\left(\Omega\tau\right)}\right]+2a\tau,$ (4)
@@
where $m=0.25$ is the modulation depth and $\Omega=\omega_{0}/20$ is the modulation frequency, which needs to be a few times smaller than $\omega_{0}$ for this application. The linear chirp coefficient $a$ is present to compensate for the zero-loss band chirp.
If an optical field with cosine frequency modulation is used, the spacing between the sidebands is also periodically modulated. The optical field in the frequency domain can be written as
[…]
@@
Ideally, only the parts of the sideband which have negative chirp must be populated. This is achieved by amplitude modulation of the optical field $\sqrt{I(\tau)}\propto\left[1+\cos{\left(\Omega\tau\right)}\right]/2$. This ensures that the light field is strong enough for a side-band to be generated only within the selected half-periods [Fig. 4(e)]. The structured side-band contains around $10\%$ of the initial electron distribution. The negatively chirped sub-pulses are separated in time by $1/2\pi\Omega$, with chirp coefficients $\mathrm{d}E/\mathrm{d}\tau=-\hbar m\omega_{0}\Omega$, and tunable compression distance $z_{\text{com}}\approx\frac{m_{e}v^{2}\alpha}{\hbar(m\omega_{0}\Omega/2a-1)}$.
@@
After propagation for $21\,\mathrm{cm}$, a train of femtosecond pulses compressed several times below the initial duration of the pulse ($\tau_{\text{com}}\approx 6\,\mathrm{fs}$) is generated [Fig. 4(f,g)]. Considering realistic limitations, such as the finite bandwidth and the finite resolution of a standard optical pulse shaper, this application is still feasible, although minor distortions are to be expected. The temporal envelope of the side-band is expected to be shortened and the sideband modulation depth might be reduced (see Supplemental Material for details [1]).
@@
Discussion—State-of-the-art light shapers enable precise control over the spectral phase of laser pulses. Our method requires broadband optical pulses with an energy bandwidth of about 0.5 eV to allow straightforward shaping through the introduction of GDD and TOD. More complex pulse shaping—such as that needed for generating periodic pulse trains—demands a light shaper with both high spectral resolution and sufficient bandwidth. For instance, a shaper with 100 pixels and a bandwidth of 0.7 eV centered around 1.2 eV is already expected to produce a pulse sufficient for the application. For further details, see Supplementary Material, Section II. [1].
@@
We need to consider the phase-matching conditions of the chosen light-electron interaction over the broad spectrum of the optical pulse. In the presented simulation results, we approximate the coupling parameter $g(\omega)$ as constant over the bandwidth of the light pulse (see derivation in SM, Section I. [1]). In reality, the spectral dependence of $g(\omega)$ must be taken into account, and it may be necessary to adjust $I(\omega)$ to achieve the desired temporal probability profile of the generated sidebands. Furthermore, we do not address the effects of transverse momentum transfer during the interaction, which are expected to introduce additional nonlinear chirp due to the unequal trajectory lengths associated with different energies within the sidebands. However, we note that this effect can be mitigated by iteratively shaping the optical field to also compensate for electron delays arising from transverse effects.
Additional temporal broadening can arise from the fact that, in a focused beam, the length of the electron trajectory depends on the propagation angle. This effect can be compensated for by spatio-temporal shaping of the envelope of the driving laser pulse, e.g., by introducing a spherical aberration to the beam leading to a curvature of the envelope in the focal plane.
@@
Conclusion—We theoretically demonstrate a technique for versatile free-electron manipulation through quantum-coherent interaction with intensity- and phase-shaped optical fields. This approach effectively transfers the state of the art of light pulse shaping techniques to the pulsed electron case, enabling precise control over the electron time-energy characteristics—including spectral width, temporal duration, and both second and higher-order dispersion—in a broadly applicable manner. We demonstrate that the quantum optical electron pulse shaper may achieve electron pulse compression of electrons to few-femtosecond pulse durations without additional spectral broadening of electrons, making this approach ideally suited for ultrafast experiments requiring high time, spatial and spectral resolutions.
Acknowledgments
@@
Data Availability— The data supporting the findings of this study are openly available at [29].
Code Availability— The code used for the numerical simulations is available upon request from the corresponding author.
@@
I Theoretical considerations
$\hat{H}=\frac{1}{2m_{e}}\left(\hat{\mathbf{p}}+e\mathbf{A}\right)^{2}=\hat{H}_{0}+\hat{H}_{\text{int}},$ (S1)
where $e>0$ is the elementary charge, $m_{e}$ is the electron mass and $\hat{\mathbf{p}}$ is the electron momentum operator. The free-particle Hamiltonian $\hat{H}_{0}=\hat{\textbf{p}}^{2}/2m_{e}$ describes the evolution of the wavepacket in free space, while the interaction Hamiltonian $\hat{H}_{\text{int}}$ describes the interaction with an electromagnetic field.
I.1 Before interaction
$\tilde{\psi}(E)=\mathcal{N}_{E}e^{\left[-\frac{(E-E_{0})^{2}}{2\sigma_{E}^{2}}\right]}e^{\left[-i\frac{\alpha(E-E_{0})^{2}}{2\hbar^{2}}\right]}e^{\left[-i\frac{\beta(E-E_{0})^{3}}{6\hbar^{3}}\right]},$ (S2)
where $\mathcal{N}_{E}$ is a normalization constant, $\sigma_{E}=\delta E/\sqrt{4\ln(2)}$ the energy width, $\alpha$ is the GDD coefficient and $\beta$ is the TOD coefficient. The time-domain representation is connected to the energy representation via Fourier transform $\psi(\tau)=\mathcal{F}\left\{\tilde{\psi}(E)\right\}$.
To account for a partially coherent state, we introduce the density matrix $\rho(\tau,\tau^{*})$
$\rho(\tau,\tau^{*})=\sum_{\tau_{0}}w_{\tau_{0}}\psi(\tau-\tau_{0})\psi^{\dagger}(\tau^{*}-\tau_{0}),$ (S3)
$w_{\tau_{0}}=\mathcal{N}_{w}\exp{\left[-\frac{8\ln{2}\tau_{0}^{2}}{\tau_{\text{pe}}^{2}}\right]},$ (S4)
In phase space we represent the electron state via its Wigner function $W(E,\tau)$, derived from the density matrix $\rho(\tau,\tau^{*})$ via Wigner-Weyl transformation [5]. To model a dispersion-free state, we set $\alpha=\beta=0$ in Eq. (S2), compute the partially incoherent density matrix $\rho_{0}(t,t^{*})$ via Eq. (S3) and apply the Wigner transformation to obtain $W_{0}(E,t)$. This state has a probability density envelope FWHM duration of $\tau_{0}=(\tau_{\text{coh}}^{2}+\tau_{\text{pe}}^{2})^{1/2}$. For a dispersed state, we use nonzero $\alpha$ and $\beta$ to get $\rho_{1}(\tau,\tau^{*})$ and $W_{1}(E,\tau)$ with an extended FWHM duration $\tau_{\text{el}}=(\tau_{0}^{2}+\tau^{2}_{\text{chirp}})^{1/2}$, where $\tau_{\text{chirp}}$ accounts for dispersion induced elongation due to propagation between the photocathode and the interaction plane.
I.2 Interaction
$\phi(\textbf{r},t)=\left[-\frac{1}{\hbar}\int^{t}_{-\infty}\hat{H}_{\text{int}}(\textbf{r}+\textbf{v}_{0}(\theta-t),\theta)\,\mathrm{d}\theta\right],$ (S5)
For a monochromatic electric field with frequency $\omega$, the vector potential is expressed in terms of the electric intensity $\mathbf{A}(\mathbf{r},t)=(-ic/\omega)\mathbf{E}(\mathbf{r},t)+c.c.$. A polychromatic optical incident field with spectrum $\tilde{\mathcal{E}}(\omega)=\sqrt{I(\omega)}e^{i\varphi(\omega)}$, where $I(\omega)$ is the spectral intensity and $\varphi(\omega)$ is the spectral phase, can be expressed as a superposition of the monochromatic components
$\textbf{E}_{i}(\mathbf{r},t)=\int_{0}^{\infty}\textrm{d}\omega\,\sqrt{I(\omega)}e^{i\varphi(\omega)}e^{-i\omega(t+\mathbf{s_{i}}\cdot\mathbf{r}/c)},$ (S6)
$\textbf{E}_{i}(\mathbf{r},t)\propto\sqrt{I(\mathbf{r},t)}\exp{\left[-i\varphi(\mathbf{r},t)\right]}.$ (S7)
Expanding the phase into polynomial series of the wave argument $(t-\mathbf{s}_{i}\cdot\mathbf{r}/c)$ we obtain $\varphi(\textbf{r},t)=\omega_{0}(t-\mathbf{s}_{i}\cdot\mathbf{r}/c)+a(t-\mathbf{s}_{i}\cdot\mathbf{r}/c)^{2}+$..., where $\omega_{0}$ is the central angular frequency and $a$ is the linear chirp coefficient.
$\textbf{E}(\mathbf{r},t)=\int_{0}^{\infty}\textrm{d}\omega\,\sqrt{I(\omega)}e^{i\varphi(\omega)}\mathbf{f}_{0}(\mathbf{r},\omega)e^{-i\omega t},$ (S8)
where $\mathbf{f}_{0}(\mathbf{r},\omega)$ contains the spatial dependence of the field component on frequency $\omega$ [32, 49].
For further evaluation, we consider 1D propagation of the electrons along $z$ with momentum vector $\mathbf{p}_{0}=(0,0,p_{0})$, group velocity vector $\mathbf{v}_{0}=(0,0,v_{0})$ and we only take into account the $z$ component of the electric field $f_{0z}$. Plugging Eq. (S8) into Eq. (S5), we evaluate the evolution operator:
$U(z,t)=\exp{\left\{-i2\text{Im}\int_{0}^{\infty}\mathrm{d}\omega\,g(\omega)\sqrt{I(\omega)}e^{i\varphi(\omega)}e^{-i\omega(t-z/v_{0})}\right\}},$ (S9)
$g(x,y,\omega)=\frac{e}{\hbar\omega}\int_{-\infty}^{z}\mathrm{d}z^{\prime}\,f_{0z}(x,y,z^{\prime};\omega)e^{i\omega z^{\prime}/v_{0}}.$ (S10)
$U(\tau)=\exp{\left\{-i2|g|\sqrt{I(\tau)}\sin{[\varphi(\tau)+\arg{g}]}\right\}}.$ (S11)
For convenience, we shift all the factors to $g$, so that $I(\tau)$ is normalized to 1 at the maximum. The density matrix describing the state of the electrons after the interaction is:
$\rho_{2}(\tau,\tau^{*})=U(\tau)\rho_{1}(\tau,\tau^{*})U^{\dagger}(\tau^{*}),$ (S12)
I.3 After interaction
As a result of the interaction, sidebands with shaped time-dependent energy evolution are generated in the Wigner function $W_{2}(E,\tau)$, which are expected to be altered by dispersive propagation in vacuum for time $\theta_{\text{prop}}$. We emphasize that $\tau$ is the detection time within the electron pulse, whereas $\theta$ is the running evolution time. The evolution is described as coordinate transformation governed by the free-electron Hamiltonian $\hat{H}_{0}$ with the Hamilton equation of motion derived from its classical analogue $H_{0}$. The new coordinates after propagation time $\theta_{\text{prop}}$ are {align} τ’-τ’_0 = τ- τ_0 + [v(E)-v0]θpropv0,
E’-E_0’ = E - E_0, where $v[E(p)]=\partial E(p)/\partial p$ is the momentum-dependent group velocity. The transformed Wigner function is $W_{3}(E,\tau,\theta_{\text{prop}})=W_{2}(E^{\prime},\tau^{\prime})$.
II Feasibility considerations
@@
$\tilde{\mathcal{E}}(\omega)\propto\exp\left[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0})^{2}}{2}-i\frac{\varphi^{\prime\prime\prime}(\omega-\omega_{0})^{3}}{6}\right],$ (S13)
$\tilde{\mathcal{E}}(\omega)\propto\sum_{k=-1}^{k=1}\sum_{n=-\infty}^{\infty}(1/2)^{|k|}i^{n}J_{n}\left(\frac{m\omega_{0}}{\Omega}\right)\times\exp\left[-\frac{(\omega-\omega_{0,kn})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0,kn})^{2}}{2}\right].$ (S14)
@@
To model the realistic light pulse, we first limit the spectral width by a rectangular window, which is non-zero for the interval $(\omega_{0}-\Delta\omega/2;\omega_{0}+\Delta\omega/2)$. Then we divide this interval by the number of pixels $N$ and assign a discrete amplitude and phase value to each of the frequencies, effectively truncating and down-sampling the light pulse frequency spectrum and spectral phase. Formally, the continuous integration in Eq. (S6) is replaced by the discrete sum
@@
On the other hand, in the case of the periodic frequency and intensity modulation of the optical pulse, the spectrum is more complicated, consisting of many peaks (see Fig. S3, blue line). While the discretization to 100 pixels is not detrimental to the peak resolution, important features of the spectrum are not transmitted through the spectral window, resulting in changes in the spectrum of the optical pulse Fig. S3, orange line. After interaction with the realistic optical pulse intensity spectrum, the sub-pulses have narrower spectrum [see Fig. S4]. However, our data demonstrate that downsampling of the spectral intensity and phase of the optical pulse shaper are not detrimental for the operation of the proposed electron pulse shaper.